# Super-Linear: A Lightweight Pretrained Mixture of Linear Experts for Time Series Forecasting

### v1

@@ 3 Super-Linear Mixture-of-Experts
4 Experiments
4.1 Experimental Details
Pre-training Datasets and Resampling.
This work emphasizes frequency diversity over sheer dataset volume. For pre-training, we use two widely adopted sources: Monash (Godahewa et al., 2021) and PEMS (Liu et al., 2022). From Monash, only datasets with more than 5,000 samples are included, ensuring sufficient data for resampling. Resampling, both upsampling and downsampling, is performed on-the-fly during training via linear interpolation, eliminating the need to store multiple copies and saving space. We use the frequencies detailed in App. C.3. The same pre-training data is used across both training stages, as outlined in Sec. 3.
Training and Implementation.
All Super-Linear models are trained on a single NVIDIA RTX3090 (24GB) using PyTorch (Paszke et al., 2019). In Stage 1, linear experts are trained independently and later combined into the Super-Linear architecture. We use the Adam optimizer with MSE loss and without auxiliary loss factors. For full-shot, the top-k experts are chosen via validation loss. All Super-Linear models use a lookback window of 512 and forecast a horizon of 96; longer horizons are predicted autoregressively. Full hyperparameters are detailed in App. E.
Baselines and Evaluations.
We compare our method to top-performing models. For zero-shot, we benchmark against Timer-XL, ROSE, Time-MoE, Moirai, TimesFM, and Chronos (Liu et al., 2024c; Wang et al., ; Shi et al., 2024; Woo et al., 2024; Das et al., 2024; Ansari et al., 2024), and focus mostly on models available on GitHub or Hugging Face. Evaluation follows the LTSF benchmark in both zero-shot and full-shot setups, using ETTm1, ETTm2, ETTh1, ETTh2, Electricity, Traffic, and Weather datasets (Liu et al., 2024c; Shi et al., 2024), including the GIFT-Eval benchmark (Aksu et al., 2024), which covers a wide range of dataset domains and sampling rates. The appendix includes additional evaluations such as full-shot (see Tab. 4), and extra model comparisons.
Figure 5: GIFT-Eval performance and parameter count of Super-Linear compared to prominent foundation models in TSF (Creation Date: 01/08/2025). The MASE score represents the geometric mean MASE across datasets, normalized by the seasonal-naive.
Figure 6: Dataset pre-training distribution for sampling rate or frequency. It is shown that other models rely heavily on hourly-sampled data. The frequencies for Super-Linear are given in App. 12.
4.2 Main Results
@@
In this section, we present the zero-shot (ZS) performance of Super-Linear in comparison to state-of-the-art (SOTA) models, given in Tab. 1. Despite having a relatively small number of parameters, Super-Linear achieves average mean squared error (MSE) reductions of 26.2%, 14.2%, 8.5%, 7.1%, and 6.3% compared to Chronos, TimesFM, Moirai, Time-MoE Large, and Time-MoE Base, respectively. Additionally, Super-Linear outperforms Timer-XL by achieving 13 best scores in MSE and 16 best scores in mean absolute error (MAE), compared to Timer-XL’s 4 and 0 best scores. Remarkably, Super-Linear accomplishes this while being only 3% the size of Timer-XL, with only 2.5M parameters compared to 84M. To ensure fairness, we compare against models with similar lookback windows, excluding evaluations that rely on excessively long lookbacks, parameter search, or ensembles.
GIFT-Eval: Model Performance and Model Size.
GIFT-Eval (Aksu et al., 2024) is a comprehensive benchmark for time series forecasting, designed to fairly assess the zero-shot and universal forecasting capabilities of foundation models across diverse domains, frequencies, prediction lengths, and variates. We evaluate Super-Linear against state-of-the-art models including Sundial, VisionTS (Liu et al., 2025; Chen et al., 2024), Chronos, and TimesFM (Ansari et al., 2024; Das et al., 2024) in Fig. 5. Despite having the smallest parameter count (2.5M), Super-Linear achieves competitive results, delivering a 16% MASE reduction over TTM, the next smallest model, while being competitive or surpassing larger Transformer-based models such as VisionTS, TimesFM, and Sundial. Note that the GIFT-Eval leaderboard is continuously updated, and our results reflect peer-reviewed models available at the time stated in Fig. 5.
Table 2: Zero-shot performance with varying sampling rates. The notations m, h and D represent per minute, hour, and day, respectively.
Model ETT1-20m ETT2-20m ETT1-30m ETT2-30m ETT1-45m ETT2-45m ETT1-12h ETT2-12h ETT1-D ETT2-D Electricity-D
Super-Linear 0.371 0.170 0.408 0.189 0.422 0.207 0.665 0.509 0.746 0.586 0.644
Timer-XL 0.841 0.199 0.437 0.208 0.747 0.239 0.705 0.772 0.811 0.536 0.945
Time-MoE 0.663 0.223 0.449 0.222 0.575 0.268 0.678 0.542 0.756 0.557 0.751
Moirai 1.013 0.208 0.516 0.211 0.893 0.254 1.211 0.691 0.914 0.722 0.834
TimesFM 0.514 0.204 0.493 0.217 0.552 0.236 0.740 0.456 0.888 1.023 0.719
Chronos 0.764 0.212 0.585 0.211 0.738 0.250 1.032 0.483 0.945 0.678 0.751
@@
Existing pre-trained models exhibit a strong reliance on hourly-sampled datasets, which comprise between 59.4% and 93.1% of their training data (Fig. 6). In contrast, Super-Linear promotes frequency diversity by training on a balanced set of predefined sampling rates using linear interpolation. To assess robustness to underrepresented frequencies, an area often overlooked in TSF, we evaluate performance across a range of resampled datasets (Tab. 2). Super-Linear achieves the best results in 9 out of 11 scenarios and demonstrates the most consistent performance overall, which we attribute to its frequency-aware training and linear expert design. To ensure fairness, we apply only upsampling (avoiding external data leakage) and follow the standard LTSF training, validation, and test protocol used for ETT and Electricity.
@@
4.3 Expert Analysis
Figure 7: Top-k expert distribution for different datasets using the Super-Linear pretrained model.
Table 3: Ablation study on the full Super-Linear setup by systematically excluding each component of the framework: long lookback resampling, complementary experts, main and naive experts, and the two-stage training scheme.
@@ 4.3 Expert Analysis
Figure 8: Top-k expert selection comparison. Performance across varying top-$k$ expert settings. Blue and red vertical lines indicate the $k$ values used during pretraining and inference, respectively.
@@ 5 Theoretical Analysis
We provide a theoretical analysis of the gating mechanism through the lens of the bias-complexity tradeoff (Shalev-Shwartz & Ben-David, 2014), yielding bounds on the frequency approximation error. This framework helps assess the strengths and limitations of our frequency-specialized approach. Specifically, it decomposes the error of the ERM algorithm into a frequency-focused approximation error $\mathcal{E}_{\text{app}}$, representing bias from limited model capacity, and estimation error $\mathcal{E}_{\text{est}}$, arising from overfitting due to model complexity. By analyzing this tradeoff, we obtain theoretical insights into the robustness of the model and the gating mechanism. Assuming $X$ and $Y(X)$ share the same periodicity, formally: $I^{X}=I^{Y}=I$, and for some $\gamma>0$ we then get,
@@
To ease notation, $G$, $\beta$, $E_{\perp}$ represent $G(X)$, $\beta(X)$,and $E_{\perp}(X)$, respectively. Adding more frequency and complementary experts reduces approximation error but risks overfitting, while too few experts lead to underfitting, reflecting the bias–complexity trade-off (Shalev-Shwartz & Ben-David, 2014). Incorporating natural frequencies C.3 introduces a meaningful inductive bias, ensuring sufficient representative experts to cover diverse datasets while controlling complexity to prevent overfitting. In A.7, we provide the proof and provide an analysis with a stricter assumption on the shared periodicity.
@@ 6 Conclusion
Appendix A Supplementary Experiments and Results
A.1 Full-Shot Forecasting
Full-Shot Forecasting.
For full-shot, we evaluate against Timer-XL, UniTST, CycleNet, iTransformer, PatchTST, GPT4TS, TimesNet, and Autoformer (Liu et al., 2024c, a; Lin et al., 2024a; Liu et al., 2023; Nie et al., 2023; Zhou et al., 2023; Wu et al., 2022, 2021). Comparisons to additional MoE-based models are given in App. 4. We evaluate Super-Linear against leading in-domain TSF models, with detailed results provided in Tab. 4. Results for CycleNet, GPT4TS, and PatchTST are drawn from their original papers, while others follow the evaluation setup of (Liu et al., 2024c). All baselines are assessed using comparable lookback windows: $L=512$ for PatchTST and GPT4TS, $L=720$ for CycleNet, and $L=672$ for the remaining models. Super-Linear uses a fixed lookback of $L=512$ and selects the best-performing model based on validation MSE across top-$k$ experts, where $k\in{6,8,10,12,20}$, with all other hyperparameters unchanged. It achieves the best MSE and MAE in 15 and 16 benchmarks, respectively—outperforming all competitors. On average, Super-Linear yields a 2.5% and 3.9% reduction in MSE compared to Timer-XL and CycleNet, respectively. Notably, it requires only a single training per dataset, with longer horizons handled autoregressively. Further, Stage 2 is repeated across three random seeds while training only on the target dataset and not on the pre-trained data.
@@ A.1 Full-Shot Forecasting
A.2 MoE and Linear Model Full-Shot Comparison
@@
A.3 Super-Linear comprehensive ZS Comparison to TTM
@@
A.4 Complementary Experts Ablation
@@
A.5 Full-Shot Results With Standard Deviation
@@
A.6 Full Ablation Results
@@
A.7 Varying Input Lookbacks
@@
1:Input: $X\in\mathbb{R}^{L}$, scales $S=\{s_{1},s_{2},\ldots\}$, energy loss penalty scaler $\lambda$, a pre-trained Super-Linear model, and a parameter $E_{\max\ \text{loss}}$ for the maximal energy loss constraint.
@@
6:for each $s\in S,s>1$ do
7:  $\omega_{\text{Nyq}}\leftarrow\tfrac{0.5}{s}$
8:  $I_{s}\leftarrow\{I_{\text{norm}}^{j}\mid\omega_{j}>\omega_{\text{Nyq}}\}$
9:  $E_{\text{lost}}\leftarrow\|I_{s}\|_{1}$
10:  $A_{\text{penalty}}[s]\leftarrow\tfrac{E_{\text{lost}}}{E_{\text{total}}+\varepsilon}$  where $\varepsilon\approx 10^{-10}$
@@
15:for each $s\in S,s>1$ do
16:  $X_{\text{interp}}\leftarrow interpolate(X,\tfrac{1}{s})\in\mathbb{R}^{L_{\text{interp}}}$
17:  if $L_{\text{train}}\leq L_{\text{interp}}$ and $s\leq s_{\max}$ then
18:   $score\leftarrow entropy(G(X_{\text{interp}}))+\lambda A_{\text{penalty}}[s]$
19:   if $score<score_{\text{best}}$ then
20:    $s_{\text{output}}\leftarrow s$
21:    $X_{\text{output}}\leftarrow X_{\text{interp}}$
@@
25:Output: $X_{\text{output}}\in\mathbb{R}^{L_{\text{interp}}},\;s_{\text{output}}$
Appendix B Theoretical Analysis
Following the bias-complexity tradeoff analysis (Shalev-Shwartz & Ben-David, 2014), we can decompose the error into the approximation error $\mathcal{E}_{\text{app}}$, and estimation error $\mathcal{E}_{\text{est}}$. The term $\mathcal{E}_{\text{app}}$ represents the minimum achievable error by the Super-Linear predictor, whereas the term $\mathcal{E}_{\text{est}}$ measures the difference between the approximation error and the error achieved by the ERM predictor with the given training data.

### v2

@@ 4.2 Main Results
GIFT-Eval: Model Performance and Model Size.
GIFT-Eval (Aksu et al., 2024) is a comprehensive benchmark for time series forecasting, designed to fairly assess the zero-shot and universal forecasting capabilities of foundation models across diverse domains, frequencies, prediction lengths, and variates. We evaluate Super-Linear against state-of-the-art models including Sundial, VisionTS Liu et al. (2025); Chen et al. (2024), Chronos, and TimesFM Ansari et al. (2024); Das et al. (2024) in Fig. 5. Despite having the smallest parameter count (2.5M), Super-Linear achieves competitive results, delivering a 16% MASE reduction over TTM, the next smallest model, while being competitive or surpassing larger Transformer-based models such as VisionTS, TimesFM, and Sundial. Note that the GIFT-Eval leaderboard is continuously updated, and our results reflect peer-reviewed models available at the time stated in Fig. 5.
@@
In this section, we present the zero-shot (ZS) performance of Super-Linear in comparison to state-of-the-art (SOTA) models, given in Tab. 1. Despite having a relatively small number of parameters, Super-Linear achieves average mean squared error (MSE) reductions of 26.2%, 14.2%, 8.5%, 7.1%, and 6.3% compared to Chronos, TimesFM, Moirai, Time-MoE Large, and Time-MoE Base, respectively. Additionally, Super-Linear outperforms Timer-XL by achieving 17 best scores in MSE and 19 best scores in mean absolute error (MAE), compared to Timer-XL’s 4 and 0 best scores. Remarkably, Super-Linear accomplishes this while being only 3% the size of Timer-XL, with only 2.5M parameters compared to 84M. To ensure fairness, we compare against models with similar lookback windows, excluding evaluations that rely on excessively long lookbacks, parameter search, or ensembles.
@@
Existing pre-trained models exhibit a strong reliance on hourly-sampled datasets, which comprise between 59.4% and 93.1% of their training data (Fig. 6). In contrast, Super-Linear promotes frequency diversity by training on a balanced set of predefined sampling rates using linear interpolation. To assess robustness to underrepresented frequencies, an area often overlooked in TSF, we evaluate performance across a range of resampled datasets (Tab. 2). Super-Linear achieves the best results in 8 out of 10 scenarios and demonstrates the most consistent performance overall, which we attribute to its frequency-aware training and linear expert design. In this setting, we follow the standard LTSF training, validation, and test protocol used for ETTm. We resample ETTm to match the corresponding sampling rate of 10, 20, 30, 40, and 50 minutes.
Figure 6: Dataset pre-training distribution for sampling rate or frequency. It is shown that other models rely heavily on hourly-sampled data. The frequencies for Super-Linear are given in App. 11.
@@
Table 2: Right: Zero-shot performance with varying sampling rates. Left: Ablation study of Super-Linear by removing key components: long lookback resampling, complementary experts, expert types, and two-stage training.
10min 20min 30min 40min 50min
Model ETTm1 ETTm2 ETTm1 ETTm2 ETTm1 ETTm2 ETTm1 ETTm2 ETTm1 ETTm2
Super-Linear 0.361 0.147 0.371 0.170 0.408 0.189 0.416 0.198 0.839 0.255
TimesFM-2.0 0.443 0.155 0.503 0.198 0.447 0.201 0.562 0.222 0.589 0.239
TimeMoE 0.616 0.197 0.663 0.223 0.449 0.220 0.618 0.269 0.657 0.281
TiRex 0.544 0.168 0.453 0.182 0.429 0.194 0.464 0.216 0.463 0.231
Moirai 1.226 0.212 0.991 0.231 1.027 0.241 1.060 0.261 1.013 0.269
Timer-XL 0.609 0.161 0.841 0.199 0.437 0.208 0.868 0.234 0.865 0.252
@@
Figure 7: Top-$k$ expert behavior in the Super-Linear model. Left: Performance comparison across varying top-$k$ expert settings, with blue and red vertical lines indicating the $k$ values used during pretraining and inference, respectively. Right: Distribution of selected experts across different datasets using the pretrained model.
@@
4.3 Inference Time
In this section, we extend the discussion and summarize the results presented in Figure 1. Across models drawn from both benchmarks, Table 3 demonstrates that Super-Linear achieves substantially faster inference, delivering $8.3\times$ speedup over Chronos-Bolt-small and $383.3\times$ over TimeMoE, while maintaining competitive accuracy. These results highlight that Super-Linear provides strong inference time performance while maintaining competitive accuracy, despite its very small parameter count. The inference time represents the average time, in seconds, required to process a batch of size 64, lookback 512, and horizon 32.
Table 3: Inference time comparison of prominent TSFMs. Each result represents the average inference time with lookback 512, horizon 32, and batch size 64.
Model Size (M) Time (sec) Speedup
Super-Linear 2.5 0.0018 $\times 1$
TTM-R2-Zeroshot 4.2 0.0053 $\times 2.9$
Timer-XL 84 0.0070 $\times 3.9$
Chronos-Bolt-Small 48 0.0150 $\times 8.3$
Chronos-Bolt-Base 205 0.0280 $\times 15.5$
Moirai-base 91 0.0380 $\times 21.1$
TimesFM 200 0.0380 $\times 21.1$
TimesFM-2.0 500 0.0630 $\times 35.0$
TiRex 35 0.0800 $\times 44.4$
VisionTS 114 0.1000 $\times 55.5$
Sundial 128 0.4400 $\times 244.4$
TimeMoE 113 0.6900 $\times 383.3$
Chronos-base 201 4.6800 $\times 2600.0$
@@ 5 Theoretical Analysis
We provide a theoretical analysis of the gating mechanism through the lens of the bias-complexity tradeoff (Shalev-Shwartz and Ben-David, 2014), yielding bounds on the frequency approximation error. This framework helps assess the strengths and limitations of our frequency-specialized approach. Specifically, it decomposes the error of the ERM algorithm into a frequency-focused approximation error $\mathcal{E}_{\text{app}}$, representing bias from limited model capacity, and estimation error $\mathcal{E}_{\text{est}}$, arising from overfitting due to model complexity. By analyzing this tradeoff, we obtain theoretical insights into the robustness of the model and the gating mechanism. Assuming $X$ and $Y(X)$ share the same periodicity, formally: $I^{X}=I^{Y}=I$, and for some $\gamma>0$ we then get,
@@
To ease notation, $G$, $\beta$, $E_{\perp}$ represent $G(X)$, $\beta(X)$,and $E_{\perp}(X)$, respectively. Adding more frequency and complementary experts reduces approximation error but risks overfitting, while too few experts lead to underfitting, reflecting the bias–complexity trade-off (Shalev-Shwartz and Ben-David, 2014). Incorporating natural frequencies D.3 introduces a meaningful inductive bias, ensuring sufficient representative experts to cover diverse datasets while controlling complexity to prevent overfitting. To further validate this analysis, in App. 7, we empirically validate our inductive bias premise, showing that selecting frequencies at random degrades performance due to increased approximation error. In B.9, we provide the proof and provide an analysis with a stricter assumption on the shared periodicity.
@@ 6 Conclusion
Appendix A Appendix
This appendix provides comprehensive supplementary material to support the main findings of our work. We present additional experimental results, ablation studies, theoretical analyses, and detailed descriptions of datasets and benchmark models. Section A covers full-shot and zero-shot forecasting results, including extensive comparisons with state-of-the-art baselines and ablation studies on model components and input lookback lengths. Section B offers a theoretical analysis of the Super-Linear model’s error decomposition. Section C details the datasets used for both pre-training and evaluation, while Section D describes the benchmark models considered in our comparisons. Finally, we provide further insights into the frequency expert weights and hyperparameter configurations used in our experiments.
@@ Appendix A Appendix
Appendix B Supplementary Experiments and Results
B.1 Full-Shot Forecasting
Full-Shot Forecasting.
For full-shot, we evaluate against Timer-XL, UniTST, CycleNet, iTransformer, PatchTST, GPT4TS, TimesNet, and Autoformer (Liu et al., 2024c; a; Lin et al., 2024a; Liu et al., 2023; Nie et al., 2023; Zhou et al., 2023; Wu et al., 2022; 2021). Comparisons to additional MoE-based models are given in App. B.1. We evaluate Super-Linear against leading in-domain TSF models, with detailed results provided in Tab. 4. Results for CycleNet, GPT4TS, and PatchTST are drawn from their original papers, while others follow the evaluation setup of Liu et al. (2024c). All baselines are assessed using comparable lookback windows: $L=512$ for PatchTST and GPT4TS, $L=720$ for CycleNet, and $L=672$ for the remaining models. Super-Linear uses a fixed lookback of $L=512$ and selects the best-performing model based on validation MSE across top-$k$ experts, where $k\in{6,8,10,12,20}$, with all other hyperparameters unchanged. It achieves the best MSE and MAE in 15 and 16 benchmarks, respectively—outperforming all competitors. On average, Super-Linear yields a 2.5% and 3.9% reduction in MSE compared to Timer-XL and CycleNet, respectively. Notably, it requires only a single training per dataset, with longer horizons handled autoregressively. Further, Stage 2 is repeated across three random seeds while training only on the target dataset and not on the pre-trained data.
B.2 MoE and Linear Model Full-Shot Comparison
@@
B.3 Super-Linear comprehensive ZS Comparison to TTM
@@
B.4 Complementary Experts Ablation
@@
Table 7: Comparison of different frequency selection strategies. Random: Select random frequencies. Uniform: select equally spaced frequencies, and ours: Select natural and common frequencies as discussed in D.3.
Ours Random Uniform
ETTm1 0.317 0.352 0.369
ETTm2 0.179 0.188 0.194
ETTh2 0.279 0.283 0.301
ETTh1 0.369 0.366 0.417
Traffic 0.414 0.424 0.477
Weather 0.159 0.165 0.161
Electricity 0.141 0.152 0.182
B.5 Frequency Choice
In this section, we compare three frequency selection strategies: uniformly spaced frequencies, randomly selected frequencies, and our method. The results are given in Table 7. Both the uniform and random baselines select 37 frequencies from the range $\tfrac{1}{512},\tfrac{1}{4}]$. We chose this interval because it contains many common natural frequencies (e.g., hourly, daily, monthly), giving these baselines a slight advantage. The rest of the training pipeline remains unchanged to ensure a fair comparison.
The results clearly show that selecting natural, commonly occurring frequencies, as done in our approach, yields substantially more consistent performance. This aligns with the theoretical bias–complexity tradeoff discussed in Section 3. The large errors observed in the uniform and random settings stem from frequency approximation error, which increases when the signal’s energy lies outside an expert’s frequency range.
B.6 Full-Shot Results With Standard Deviation
@@
B.7 Full Ablation Results
@@
B.8 Varying Sampling Rates
Extending the discussion from Sec. 4.2 on zero-shot generalization across varying sampling rates, we include a comparison on the GIFT-Eval benchmark using the BizItObs dataset at a 10-second sampling rate. This sampling rate is rarely present in the training corpora of the evaluated foundation models, and for Super-Linear, it also falls outside the range of its frequency experts. Thus, this dataset serves as an additional benchmark for evaluating generalization to unseen frequencies. The results show that Super-Linear achieves an average score of 2.05, slightly lower than TiREX and Sundial, but substantially better than Chronos-Bolt ($\sim$3) and Timer ($\sim$2.5), highlighting the strong generalization capability of Super-Linear on uncommon sampling rates.
Table 10: We compare the performance of Super-Linear on datasets with sampling frequencies outside its training regime, specifically at the 10-second sampling rate on the BizItObs dataset from GIFT-Eval.
dataset TiRex Sundial Super-Linear Timer Chronos Base TTM R2 Chronos Bolt Small Moirai Base TimesFM
bizitobs application 10S long 3.40 3.71 3.81 8.51 9.25 9.59 9.65 13.50 16.30
bizitobs application 10S medium 2.64 2.86 3.06 7.47 9.87 9.02 9.15 12.80 11.20
bizitobs application 10S short 1.21 1.43 1.55 3.75 3.01 4.21 5.41 5.32 4.36
bizitobs service 10S long 1.49 1.46 1.57 4.25 4.22 5.34 4.85 6.08 7.77
bizitobs service 10S medium 1.17 1.27 1.39 4.26 4.58 5.12 4.64 5.99 6.46
bizitobs service 10S short 0.82 0.84 0.94 2.17 1.88 2.73 2.90 3.43 2.90
Average 1.79 1.93 2.05 5.07 5.47 6.00 6.10 7.85 8.16
B.9 Varying Input Lookbacks
@@
1:Input $X\in\mathbb{R}^{L}$, scales $S=\{s_{1},s_{2},\ldots\}$, energy loss penalty scaler $\lambda$, a pre-trained Super-Linear model, and a $E_{max\ loss}$ parameter to act as a hard maximal energy loss constrain.
@@
6:for $s\in S,s>1$ do
7:  $\omega_{\text{Nyq}}\leftarrow\frac{0.5}{s}$, marks the threshold from which higher frequencies will be lost.
8:  $I_{s}\leftarrow\left\{I_{norm}^{j}\;\middle|\;\omega_{j}>\omega_{\text{Nyq}}\right\}\in\mathbb{R}^{<M},$
[…]
10:  $A_{penalty}[s]\leftarrow\frac{E_{\text{lost}}}{E_{\text{total}}+\varepsilon},\quad\varepsilon\approx 10^{-10}$, the fraction of normalized energy loss to be used as a penalty.
@@
15:for $s\in S,s>1$ do
16:  $X_{interp}\leftarrow interpolate(X,\frac{1}{s})\in\mathbb{R}^{L_{interp}}$
17:  if $L_{train}\leq L_{interp}$ and $s\leq s_{max}$ then
18:   $s\leftarrow entropy(G(X_{interp}))\ +\lambda A_{penalty}[s]$
19:   if $score<score_{best}$ then
20:     $s_{output}\leftarrow s$
21:     $X_{output}\leftarrow X_{interp}$
@@
25:return $X_{output}\in\mathbb{R}^{L_{interp}},s_{output}$
Appendix C Theoretical Analysis
Following the bias-complexity tradeoff analysis (Shalev-Shwartz and Ben-David, 2014), we can decompose the error into the approximation error $\mathcal{E}_{\text{app}}$, and estimation error $\mathcal{E}_{\text{est}}$. The term $\mathcal{E}_{\text{app}}$ represents the minimum achievable error by the Super-Linear predictor, whereas the term $\mathcal{E}_{\text{est}}$ measures the difference between the approximation error and the error achieved by the ERM predictor with the given training data.